Dear tutor I really need ur help I newly moved to US and did not learn some topics and need help asap and would be so thankful for ur help. Here is my question 11ml of a saline solution was mixed with 4ml of a 55% saline solution to make a 77% saline solution. Find the percent concentration of the first solution.
Answer by ankor@dixie-net.com(22740)

11ml of a saline solution was mixed with 4ml of a 55% saline solution to make a 77% saline solution.
Find the percent concentration of the first solution.
:
Let x = the percent concentration in decimal form
We know the resulting amt is 11 + 4 = 15 ml
:
11x + .55(4) = .77(15)
11x + 2.2 + = 11.55
11x = 11.55 - 2.2
11x = 9.35
x = 9.35/11
x = .85
:
the first solution was an 85% solution
:
;
Check this
.85(11) + .55(4) = .77(15)